Find cosθ if sinθ = 3/5 and 90° < θ <180°.
Thank You

Answer by fractalier(2101)

Let q be theta...
Since sin^2 q + cos^2 q = 1, we know that
cos^2 q = 1 - sin^2 q
cos^2 q = 1 - (3/5)^2
cos^2 q = 1 - 9/25 = 16/25 and then
cos q = 4/5
But in the 2nd quadrant, cos q is negative so the answer is
-4/5